OpenStudy (angelberina):

What is the derivative of h(t) = ln(7)7^t

zepdrix (zepdrix):

\[\large\rm h(t)=(\ln7)7^t\]Derivative of this guy?

zepdrix (zepdrix):

You might be tempted to use product rule, but don't! ln7 contains no t. It's just a constant. It might look like something more, but it's just a number dressed up in a fancy tuxedo.

zepdrix (zepdrix):

So ignore that while differentiating. Remember how to differentiate an exponential? Here is our rule,\[\large\rm \frac{d}{dx}a^x=a^x(\ln a)\]

zepdrix (zepdrix):

So then,\[\large\rm \frac{d}{dt}7^t=?\]What do you think? :d

OpenStudy (angelberina):

Thank you!